# The Gray Zone of Faithfulness: Taming Ambiguity in Unfaithfulness Detection

### v2

@@ 1 Introduction
Knowledge-grounded generation, such as summarization, extends LLMs’ application to domain-specific tasks, where faithfulness to the provided source knowledge is critical, especially in high-stakes fields like finance and healthcare. However, SOTA LLMs can still generate content unfaithful to given knowledge sources, which yields the faithfulness hallucination (or unfaithfulness for brevity) detection task (Niu et al., 2024; Cossio, 2025).
Unlike general hallucination detection, unfaithfulness detection specifically identifies errors unfaithful to a given knowledge source, e.g., a document. To evaluate unfaithfulness detectors, various benchmarks have been proposed for tasks including knowledge question answering (Dziri et al., 2022; Liu et al., 2023; Sadat et al., 2023; Niu et al., 2024; Ji et al., 2024) and summarization (Bao et al., 2025). Despite these advances, existing benchmarks face a fundamental challenge: annotation ambiguity due to the ill-defined boundary of permissible common sense in model outputs, as noted by Seo et al. (2025). The ambiguity leads to inconsistent evaluations, undermining the reliability of faithfulness assessments.
@@
Specifically, current unfaithfulness detection benchmarks do not consider the annotators’ divergent understandings of the model outputs. For instance, the widely adopted AIS framework (Rashkin et al., 2023) relies on an idealized “generic hearer,” yet fails to account for annotators’ varying cultural backgrounds and domain expertise. As also reported by Seo et al. (2025), we observe that such ambiguity manifests in existing benchmarks – for instance, a golf-related claim that interprets “the green jacket” as “the Masters Tournament winner” (see Table 1) can be considered fully supported only by those familiar with golf. While FaithBench (Bao et al., 2025) attempts to address this by introducing categories Benign and Questionable for ambiguous cases, their definitions remain vague and do not ensure annotation quality (see Sec. 2 for more details).
To overcome these limitations, we propose a novel annotation framework for LLM faithfulness that clearly labels the ambiguity. Our key innovation is the rigorous definitions of two intermediate categories, Out-Dependent and Ambiguous, to capture cases where verification depends on external knowledge and on different interpretations. These definitions offer an efficient alternative to extensive crowdsourcing efforts, such as those employed by Glockner et al. (2024). Using this framework, we construct VeriGray, a novel unfaithfulness detection benchmark for summarization, enabling more objective and granular evaluations. Analysis of our benchmark reveals that even SOTA LLMs are prone to hallucination on the summarization task, and a substantial proportion of generated sentences fall into the Out-Dependent category, underscoring the importance of resolving annotation ambiguity in unfaithfulness detection benchmarks. Moreover, evaluations on VeriGray show that current detection methods face significant challenges, especially in identifying Out-Dependent and Ambiguous cases, pointing to substantial room for future improvement.
@@
Cao and Wang (2021) BART, PEGASUS span-level experts
BEGIN (Dziri et al., 2022) T5-base, GPT2-base response-level human
VERI-GRAN (Liu et al., 2023) BingChat, NeevaAI, etc. citation-level crowd source
HaluEval (Li et al., 2023) ChatGPT response-level ChatGPT
DelucionQA (Sadat et al., 2023) ChatGPT sentence-level crowd source
ExpertQA Malaviya et al. (2024) GPT-4, BingChat sentence-level crowd source & experts
ANAH (Ji et al., 2024) GPT-3.5 sentence-level human assisted by GPT-4
RAGTruth (Niu et al., 2024) GPT-4, GPT-3.5, etc. word-level crowd source
AmbiFC (Glockner et al., 2024) Human sentence-level crowd source
FaithBench (Bao et al., 2025) GPT-4o, GPT-3.5, etc. span-level graduate students
@@ 2 Related Work
Unfaithfulness Detection Benchmarks. A key limitation in existing benchmarks for unfaithfulness detection is annotation ambiguity, as noted by Seo et al. (2025) (see Table 2 for a summary). Most benchmarks overlook the potential differences in cultural backgrounds and domain expertise among annotators. For instance, the widely used AIS annotation framework (Rashkin et al., 2023) assumes a “generic hearer” as the annotator. However, in practice, annotators can be diverse, and the annotation guidelines often fail to specify the extent to which common sense or external knowledge can be used to verify LLMs’ generation, leading to ambiguity in annotation. To illustrate, consider the example presented in Table 1.
Several benchmarks have attempted to address the ambiguity of unfaithfulness annotation, including the one introduced by Cao and Wang (2021), AmbiFC (Glockner et al., 2024), and FaithBench (Bao et al., 2025), but each has notable limitations. As far as we know, Cao and Wang (2021) were the first to consider ambiguity stemming from external knowledge and constructed a corresponding benchmark. Unfortunately, their work relied on old summarization models such as BART (Lewis et al., 2020) and PEGASUS (Zhang et al., 2020), and their insights have not been widely adopted in subsequent research. AmbiFC, on the other hand, focuses solely on linguistic ambiguity, not covering ambiguity induced by external knowledge. FaithBench attempts to handle ambiguity by introducing two label categories – Benign and Questionable – which refer, respectively, to cases that are “hallucination, but supported by world knowledge, common sense, or logical reasoning, such that a reader finds it acceptable or welcomed” and cases where “classification may differ depending on whom you ask”, respectively. Yet these definitions remain subjective and difficult to operationalize consistently. As shown in Table 5 of Appendix A, several annotation errors are associated with these two categories. In contrast to FaithBench, we propose a more rigorous annotation framework designed to systematically address this issue.
Automatic Ambiguity Detection. Seo et al. (2025) identify annotation errors and ambiguity in existing fact-checking benchmarks and propose an automatic detection method that leverages the discrepancies between the human annotation and predictions from multiple LLM-as-a-judge assessments. The ambiguity patterns they identify fall into four categories, which can be summarized by two categories: ambiguity related to external-knowledge-related and linguistics-related ambiguity, which align with our Out-Dependent and Ambiguous classes, respectively. However, the automatic detector cannot distinguish between annotation ambiguity and annotation errors, which can result in low precision for ambiguity detection. In this paper, we employ this ambiguity detector as part of our annotation refinement process.
@@ 3 Annotation Framework
Document Response Annotation
The “black box” of the Su-24 jet was officially opened in Moscow on Friday in front of journalists and diplomats. …The Su-24 was shot down by F-16 fighters on 24 November. … The “black box” of the Su-24 jet, shot down on November 24th, was opened in Moscow on Friday. … Explicitly-Supported
@@
…the ace was nicklaus ’ first on the hallowed turf of augusta , but the veteran saw his achievement eclipsed by that of camilo villegas on a day which saw five hole-in-ones , levelling the record from 2002 . …Despite his achievement, Camilo Villegas stole the show by scoring two holes-in-one, tying the record. Fabricated. Reason: The record in the document does not refer to the two ace of Villegas.
conor mcgregor is gearing up for a featherweight title challenge against jose aldo on july 11 …aldo makes the eighth defence of his belt against the irish fighter in las vegas , but mcgregor claimed last week that the man he is challenging lacks the same motivation as him . This is McGregor’s eighth defense of the title, and the two fighters have a contentious relationship. … Contradicting. Reason: It’s Aldso’s 8th defense.
Table 3: Examples of our taxonomy (see Table 1 for the example of Out-Dependent).
@@
No-Fact Class. As identifying verification-worthy sentences is a prerequisite of checking the facts in the generated text (Liu et al., 2023), we first categorize non-verification-worthy sentences as the No-Fact class. This category includes sentences that do not convey any factual content, e.g., “would you like to learn more?”.
Faithful Classes. We observe that a large amount of the LLM-generated summary sentences are paraphrases of the original document, making them straightforward to verify. Motivated by selective prediction (Chow, 1957; Geifman and El-Yaniv, 2017), we argue that this presents an opportunity for the automatic detectors to accurately classify such easy cases, leaving the more challenging minority of samples to be verified by human experts. This approach can help reduce the cost of manual fact-checking. Additionally, annotators can assign a special label to these sentences to indicate their high confidence in their faithfulness. Accordingly, we classify such sentences as Explicitly-Supported and leave the remaining supported sentences to Implicitly-Supported. We formalize the definitions as follows.
@@
Unfaithful Classes. According to the severity of unfaithfulness, we divide unfaithfulness into two classes: Contradicting and Fabricated, which correspond respectively to the Contradiction and Neutral classes of in Natural Language Inference (NLI) task (Nie et al., 2020). Contradicting aligns directly with the standard Contradiction category in NLI. However, we define Fabricated with a key distinction from Neutral: it excludes sentences that are supported by external knowledge, as such sentences are instead classified as Out-Dependent (discussed later). The formal definitions are as follows.
@@ 4 VeriGray: A dataset of Objective Unfaithfulness Annotation
Considering that summarization is a typical scenario for unfaithfulness detection (Scialom et al., 2021; Pagnoni et al., 2021; Niu et al., 2024; Bao et al., 2025), we build a summarization dataset annotated with unfaithfulness labels, as detailed below.
@@ 4.1 Data Collection
We collected the documents from FaithBench (Bao et al., 2025), whose passages for summarization come from various NLI, fact-checking, and summarization datasets. To exclude NLI instances, we removed instances where the document is shorter than twice the length of its corresponding summary. We also removed self-contradictory documents. The remaining documents, along with their LLM-generated summaries, were retained, but original faithfulness annotations were removed. Subsequently, we employed more recent LLMs, including GPT-5 OpenAI (2025), DeepSeek V3-0324 (DeepSeek-AI, 2024), and Qwen3-8B (Yang et al., 2025), to generate additional summaries for each document. Following the setup in FaithBench, we used the summarizer prompt of Vectara’s Hallucination Leaderboard Hughes et al. (2023). The full prompt can be found in Appendix B. We set the temperatures of DeepSeek-V3 and Qwen3-8B to 0.3 and 0.6, respectively (the temperature of GPT-5 was not adjustable via the API). After appending the newly generated summaries, our dataset comprised 412 summaries, containing a total of 2044 sentences. Each summary was segmented into sentences using NLTK (Bird, 2006), with manual corrections applied by annotators as needed.
@@ 4.2 Human Annotation
Attribution-assisted Annotation. We observed that the most time-consuming aspect of faithfulness annotation is locating relevant evidence spans in the document. To streamline this process, we integrated an attention-based fine-grained attribution method (Ding et al., 2025) into a web-based annotation tool. When an annotator selects a target sentence, the attribution module automatically highlights relevant text segments in the document, facilitating efficient evidence identification.
Quality Assurance. The quality assurance process was conducted in multiple stages to ensure high quality. Firstly, before annotation, all annotators completed a training phase that included reviewing canonical examples and passing a quiz on challenging cases. Secondly, during annotation, annotators selected not only the label but also intermediate decision options such as Contains Fast, Is Ambiguous, and Is Supported by Doc. The system automatically verified the consistency between these intermediate choices and the final label, and only consistent annotations could be submitted. Thirdly, after annotation, each instance underwent successive reviews by a second annotator. Fourthly, instances not classified as Ambiguous or Out-Dependent were processed by a modified version [^3] The original error detector runs four LLM-as-a-judge models, with its final prediction confirmed by a second round of LLM-as-a-judge. Our modification introduced a new LLM-as-a-judge model of GPT-5 and omitted the second round as it empirically reduced error detection recall compared to the first round alone. of the automatic annotation error detector proposed by Seo et al. (2025). Specifically, the modified detector runs five LLM-as-a-judge models (using o3-mini, GPT-4o, Gemini 2.0-Flash, Llama3.1 405B, and GPT-5), each providing their individual faithfulness predictions (Attributable, Not Attributable, or Contradictory). Potential annotation errors were flagged whenever any LLM’s prediction disagreed with the annotated label. Finally, error detection results were reviewed by the annotators to produce the final annotations.
Figure 2: Model-wise annotation statistics, where the models are arranged from left to right according to the descending order of the amount of unfaithful classes.
@@ 4.3 Dataset Analysis
Here, we investigate the distribution of unfaithfulness classes and the sequential dependency of unfaithfulness in our benchmark.
Distribution of Unfaithfulness Classes. Figure 2 shows the sentence-level category breakdown across different models. Overall, the majority are Faithful ($76\%$), with Explicitly-Supported sentences ($63\%$) substantially outnumbering Implicitly-Supported sentences ($13\%$). Unfaithful sentences constitute approximately $14\%$ of the total, most of which are Fabricated ($11\%$). A notable proportion of sentences fall under the Not Sure category, primarily as Out-Dependent ($8\%$), while fewer than $1\%$ are labeled as Ambiguous. Across models, DeepSeek-V3-0324 and GPT-5 yield the highest proportion of Not Sure sentences ($15\%$). Meanwhile, Claude-3.5-Sonnet generates the fewest Unfaithful sentences ($3\%$), whereas Qwen3-8B generates the most ($29\%$), reflecting the comparative success of proprietary models in maintaining faithfulness.
Sequential Dependency of Unfaithfulness. The phenomenon of hallucination snowballing (Zhang et al., 2024b) – where earlier incorrect claims trigger subsequent hallucinated explanations – has been observed in several multi-step question-answering datasets, indicating a sequential dependency of hallucinations. To examine whether such sequential dependency exists in our benchmark, we model sentence-wise faithfulness in each response as a Markov chain and estimate the transition probabilities among unfaithfulness classes (see Figure 3). As the results show, the confidence intervals of $\hat{p}(\text{Faithful}\mid\cdot)$ (and similarly for $\hat{p}(\text{Unfaithful}\mid\cdot)$) overlap across all source classes (except the outlier of the No-Fact class due to data scarcity). Comparable results are observed for model-specific and longer-dependency transition probabilities (see Appendix C for details). Thus, the sequential dependency of unfaithfulness in our dataset is weak, suggesting that prior sentence hallucinations do not provide a reliable shortcut for predicting unfaithfulness in subsequent sentences.

### v3

@@ 1 Introduction
Knowledge-grounded generation, such as summarization, extends LLMs’ application to domain-specific tasks, where faithfulness to the provided source knowledge is critical, especially in high-stakes fields like finance and healthcare. However, SOTA LLMs can still generate content unfaithful to given knowledge sources, which yields the faithfulness hallucination (or unfaithfulness for brevity) detection task (niu2024ragtruth; cossio2025comprehensive).
Unlike general hallucination detection, unfaithfulness detection specifically identifies errors unfaithful to a given knowledge source, e.g., a document. To evaluate unfaithfulness detectors, various benchmarks have been proposed for tasks including knowledge question answering (dziri2022evaluating; liu2023evaluating; sadat2023delucionqa; niu2024ragtruth; ji2024anah) and summarization (bao2025faithbench). Despite these advances, existing benchmarks face a fundamental challenge: annotation ambiguity due to the ill-defined boundary of permissible common sense in model outputs, as noted by seo2025verifying. The ambiguity leads to inconsistent evaluations, undermining the reliability of faithfulness assessments.
@@
Specifically, current unfaithfulness detection benchmarks do not consider the annotators’ divergent understandings of the model outputs. For instance, the widely adopted AIS framework (rashkin2023measuring) relies on an idealized “generic hearer,” yet fails to account for annotators’ varying cultural backgrounds and domain expertise. As also reported by seo2025verifying, we observe that such ambiguity manifests in existing benchmarks – for instance, a golf-related claim that interprets “the green jacket” as “the Masters Tournament winner” (see Table 1) can be considered fully supported only by those familiar with golf. While FaithBench (bao2025faithbench) attempts to address this by introducing categories Benign and Questionable for ambiguous cases, their definitions remain vague and do not ensure annotation quality (see Sec. 2 for more details).
To overcome these limitations, we propose a novel annotation framework for LLM faithfulness that clearly labels the ambiguity. Our key innovation is the rigorous definitions of two intermediate categories, Out-Dependent and Ambiguous, to capture cases where verification depends on external knowledge and on different interpretations. These definitions offer an efficient alternative to extensive crowdsourcing efforts, such as those employed by glockner2024ambifc. Using this framework, we construct VeriGray, a novel unfaithfulness detection benchmark for summarization, enabling more objective and granular evaluations. Analysis of our benchmark reveals that even SOTA LLMs are prone to hallucination on the summarization task, and a substantial proportion of generated sentences fall into the Out-Dependent category, underscoring the importance of resolving annotation ambiguity in unfaithfulness detection benchmarks. Moreover, evaluations on VeriGray show that current detection methods face significant challenges, especially in identifying Out-Dependent and Ambiguous cases, pointing to substantial room for future improvement.
@@
cao2021cliff BART, PEGASUS span-level experts
BEGIN (dziri2022evaluating) T5-base, GPT2-base response-level human
VERI-GRAN (liu2023evaluating) BingChat, NeevaAI, etc. citation-level crowd source
HaluEval (li2023halueval) ChatGPT response-level ChatGPT
DelucionQA (sadat2023delucionqa) ChatGPT sentence-level crowd source
ExpertQA malaviya2024expertqa GPT-4, BingChat sentence-level crowd source & experts
ANAH (ji2024anah) GPT-3.5 sentence-level human assisted by GPT-4
RAGTruth (niu2024ragtruth) GPT-4, GPT-3.5, etc. word-level crowd source
AmbiFC (glockner2024ambifc) Human sentence-level crowd source
FaithBench (bao2025faithbench) GPT-4o, GPT-3.5, etc. span-level graduate students
@@ 2 Related Work
Unfaithfulness Detection Benchmarks. A key limitation in existing benchmarks for unfaithfulness detection is annotation ambiguity, as noted by seo2025verifying (see Table 2 for a summary). Most benchmarks overlook the potential differences in cultural backgrounds and domain expertise among annotators. For instance, the widely used AIS annotation framework (rashkin2023measuring) assumes a “generic hearer” as the annotator. However, in practice, annotators can be diverse, and the annotation guidelines often fail to specify the extent to which common sense or external knowledge can be used to verify LLMs’ generation, leading to ambiguity in annotation. To illustrate, consider the example presented in Table 1.
Several benchmarks have attempted to address the ambiguity of unfaithfulness annotation, including the one introduced by cao2021cliff, AmbiFC (glockner2024ambifc), and FaithBench (bao2025faithbench), but each has notable limitations. As far as we know, cao2021cliff were the first to consider ambiguity stemming from external knowledge and constructed a corresponding benchmark. Unfortunately, their work relied on old summarization models such as BART (lewis2020bart) and PEGASUS (zhang2020pegasus), and their insights have not been widely adopted in subsequent research. AmbiFC, on the other hand, focuses solely on linguistic ambiguity, not covering ambiguity induced by external knowledge. FaithBench attempts to handle ambiguity by introducing two label categories – Benign and Questionable – which refer, respectively, to cases that are “hallucination, but supported by world knowledge, common sense, or logical reasoning, such that a reader finds it acceptable or welcomed” and cases where “classification may differ depending on whom you ask”, respectively. Yet these definitions remain subjective and difficult to operationalize consistently. As shown in Table 5 of Appendix A, several annotation errors are associated with these two categories. In contrast to FaithBench, we propose a more rigorous annotation framework designed to systematically address this issue.
Automatic Ambiguity Detection. seo2025verifying identify annotation errors and ambiguity in existing fact-checking benchmarks and propose an automatic detection method that leverages the discrepancies between the human annotation and predictions from multiple LLM-as-a-judge assessments. The ambiguity patterns they identify fall into four categories, which can be summarized by two categories: ambiguity related to external-knowledge-related and linguistics-related ambiguity, which align with our Out-Dependent and Ambiguous classes, respectively. However, the automatic detector cannot distinguish between annotation ambiguity and annotation errors, which can result in low precision for ambiguity detection. In this paper, we employ this ambiguity detector as part of our annotation refinement process.
@@ 3 Annotation Framework
Document Target Sentence Annotation
The “black box” of the Su-24 jet was officially opened in Moscow on Friday in front of journalists and diplomats. …The Su-24 was shot down by F-16 fighters on 24 November. … The “black box” of the Su-24 jet, shot down on November 24th, was opened in Moscow on Friday. Explicitly-Supported
@@
…the ace was nicklaus ’ first on the hallowed turf of augusta , but the veteran saw his achievement eclipsed by that of camilo villegas on a day which saw five hole-in-ones , levelling the record from 2002 . Despite his achievement, Camilo Villegas stole the show by scoring two holes-in-one, tying the record. Fabricated. Reason: The record in the document does not refer to the two ace of Villegas.
conor mcgregor is gearing up for a featherweight title challenge against jose aldo on july 11 …aldo makes the eighth defence of his belt against the irish fighter in las vegas , but mcgregor claimed last week that the man he is challenging lacks the same motivation as him . This is McGregor’s eighth defense of the title, and the two fighters have a contentious relationship. Contradicting. Reason: It’s Aldso’s 8th defense.
the zhangzhou plant was slated for xiamen …however , it provoked an angry backlash in 2007 due to pollution concerns and prompted the local government to relocate the factory to its current, more remote location. This plant had experienced a previous explosion in 2013 and was relocated from Xiamen due to pollution concerns and protests in 2007. Ambiguous (Explicitly-Supported / Fabricated). Reason: While the sentence is accurate when “in 2007” modifies “protests,” the alternative reading – that it specifies the time of relocation – lacks support.
Table 3: Examples of our taxonomy (see Table 1 for the example of Out-Dependent). The key segment that drives the annotation decision is in red.
@@
No-Fact Class. As identifying verification-worthy sentences is a prerequisite of checking the facts in the generated text (liu2023evaluating), we first categorize non-verification-worthy sentences as the No-Fact class. This category includes sentences that do not convey any factual content, e.g., “would you like to learn more?”.
Faithful Classes. We observe that a large amount of the LLM-generated summary sentences are paraphrases of the original document, making them straightforward to verify. Motivated by selective prediction (chow1957optimum; geifman2017selective), we argue that this presents an opportunity for the automatic detectors to accurately classify such easy cases, leaving the more challenging minority of samples to be verified by human experts. This approach can help reduce the cost of manual fact-checking. Additionally, annotators can assign a special label to these sentences to indicate their high confidence in their faithfulness. Accordingly, we classify such sentences as Explicitly-Supported and leave the remaining supported sentences to Implicitly-Supported. We formalize the definitions as follows.
@@
Unfaithful Classes. According to the severity of unfaithfulness, we divide unfaithfulness into two classes: Contradicting and Fabricated, which correspond respectively to the Contradiction and Neutral classes of in Natural Language Inference (NLI) task (nie2020adversarial). Contradicting aligns directly with the standard Contradiction category in NLI. However, we define Fabricated with a key distinction from Neutral: it excludes sentences that are supported by external knowledge, as such sentences are instead classified as Out-Dependent (discussed later). The formal definitions are as follows.
@@ 4 VeriGray: A dataset of Objective Unfaithfulness Annotation
Considering that summarization is a typical scenario for unfaithfulness detection (scialom2021questeval; pagnoni2021understanding; niu2024ragtruth; bao2025faithbench), we build a summarization dataset annotated with unfaithfulness labels, as detailed below.
@@ 4.1 Data Collection
We collected the documents from FaithBench (bao2025faithbench), whose passages for summarization come from various NLI, fact-checking, and summarization datasets. To exclude NLI instances, we removed instances where the document is shorter than twice the length of its corresponding summary. We also removed self-contradictory documents. The remaining documents, along with their LLM-generated summaries, were retained, but original faithfulness annotations were removed. Subsequently, we employed more recent LLMs, including GPT-5 openai2025introducing, DeepSeek V3-0324 (liu2024deepseek), and Qwen3-8B (yang2025qwen3), to generate additional summaries for each document. Following the setup in FaithBench, we used the summarizer prompt of Vectara’s Hallucination Leaderboard hughes2023vectara. The full prompt can be found in Appendix B. We set the temperatures of DeepSeek-V3 and Qwen3-8B to 0.3 and 0.6, respectively (the temperature of GPT-5 was not adjustable via the API). After appending the newly generated summaries, our dataset comprised 412 summaries, containing a total of 2044 sentences. Each summary was segmented into sentences using NLTK (bird2006nltk), with manual corrections applied by annotators as needed.
@@ 4.2 Human Annotation
Attribution-assisted Annotation. We observed that the most time-consuming aspect of faithfulness annotation is locating relevant evidence spans in the document. To streamline this process, we integrated an attention-based fine-grained attribution method (ding2025attention) into a web-based annotation tool. When an annotator selects a target sentence, the attribution module automatically highlights relevant text segments in the document, facilitating efficient evidence identification.
Quality Assurance. The quality assurance process was conducted in multiple stages to ensure high quality. Firstly, before annotation, all annotators completed a training phase that included reviewing canonical examples and passing a quiz on challenging cases. Secondly, during annotation, annotators selected not only the label but also intermediate decision options such as Contains Fast, Is Ambiguous, and Is Supported by Doc. The system automatically verified the consistency between these intermediate choices and the final label, and only consistent annotations could be submitted. Thirdly, after annotation, each instance underwent successive reviews by a second annotator. Fourthly, instances not classified as Ambiguous or Out-Dependent were processed by a modified version [^3] The original error detector runs four LLM-as-a-judge models, with its final prediction confirmed by a second round of LLM-as-a-judge. Our modification introduced a new LLM-as-a-judge model of GPT-5 and omitted the second round as it empirically reduced error detection recall compared to the first round alone. of the automatic annotation error detector proposed by seo2025verifying. Specifically, the modified detector runs five LLM-as-a-judge models (using o3-mini, GPT-4o, Gemini 2.0-Flash, Llama3.1 405B, and GPT-5), each providing their individual faithfulness predictions (Attributable, Not Attributable, or Contradictory). Potential annotation errors were flagged whenever any LLM’s prediction disagreed with the annotated label. Finally, error detection results were reviewed by the annotators to produce the final annotations.
Figure 2: Model-wise annotation statistics, where the models are arranged from left to right according to the descending order of the proportion of unfaithful classes.
@@ 4.3 Dataset Analysis
Here, we investigate the distribution of unfaithfulness classes in our benchmark [^4] We also explore the sequential dependency of unfaithfulness in Appendix C to see if the phenomenon of hallucination snwoballing (zhang2024language) exists in our benchmark.. Figure 2 shows the sentence-level category breakdown across different models. Overall, the majority are Faithful (76%), with Explicitly-Supported sentences (63%) substantially outnumbering Implicitly-Supported sentences (13%). Unfaithful sentences constitute approximately 14% of the total, most of which are Fabricated (11%). A notable proportion of sentences fall under the Not Sure category, primarily as Out-Dependent (8%), while fewer than 1% are labeled as Ambiguous. Across models, DeepSeek-V3-0324 and GPT-5 yield the highest proportion of Not Sure sentences (15%). Meanwhile, Claude-3.5-Sonnet generates the fewest Unfaithful sentences (3%), whereas Qwen3-8B generates the most (29%), reflecting the comparative success of proprietary models in maintaining faithfulness.